# Controllable Video Generation: A Survey

### v1

@@ 4.1.1 Pose-Guided Generation
Pose-guided video generation specifically aims to animate subjects, predominantly humans, according to a sequence of predefined poses. This is a critical task for applications such as virtual avatar animation, character generation for interactive entertainment, human motion synthesis from sparse inputs, and fashion video synthesis [76, 369, 266, 370]. The core challenge lies in generating temporally coherent video frames where the subject’s appearance is preserved while accurately following the target pose sequence $P={p_{1},p_{2},...,p_{T}}$, where each $p_{t}$ represents the pose in frame $t$.
@@
The input pose information $p_{t}$ is typically represented as 2D skeletal keypoints (e.g., COCO format), 3D skeletal coordinates, or more dense representations like DensePose [371], which maps image pixels to 3D surface coordinates of the human body. These pose sequences can be extracted from existing videos using off-the-shelf pose estimators (e.g., OpenPose [372]), derived from motion capture (MoCap) data, or generated algorithmically. The choice of pose representation influences the granularity of control and the complexity of the generation task.
@@
Numerous approaches have been proposed to achieve pose-guided generation. Early methods [78] often rely on GANs, using pose information to condition the generator, for instance, by rendering pose stick figures and feeding them as input alongside a latent code. More recent work leverages the power of diffusion models. For example, MagicAnimate [72] and Animate Anyone [71] and its successor Animate Anyone 2 [63] utilize diffusion models conditioned on pose sequences and a reference image to generate temporally consistent animations. Techniques like Champ [69] and MimicMotion [68] focus on high-fidelity human motion synthesis. DirectorLLM [64] employs large language models to direct human-centric video generation based on pose and textual descriptions. Follow Your Pose [2] introduces a two-stage training strategy to create high-quality character videos. ControlNet-based approaches [373] have also been adapted, where pose maps serve as direct spatial conditions for pre-trained text-to-image diffusion models, which are then fine-tuned or augmented with temporal modules for video tasks [67, 75]. Disentangling pose from appearance is a key strategy, often achieved by dedicated appearance encoders and pose encoders, with mechanisms like feature warping or attention to align appearance features with the target pose [66]. Temporal consistency is often enforced through temporal attention mechanisms across frames or by incorporating temporal smoothness losses.
@@
Despite significant progress, pose-guided generation still faces challenges. Handling severe self-occlusions, where parts of the subject are hidden due to the pose, requires robust inpainting capabilities. Maintaining fine-grained texture details and consistent identity across large pose variations remains difficult, especially when appearance must be reliably transferred from a reference image without explicit identity control. Ensuring natural and smooth transitions between poses, avoiding jittery or robotic movements, is crucial for realism. Furthermore, generalization to unseen poses, body shapes, or complex clothing items not well-represented in training data can be problematic, often requiring extensive and diverse datasets like those proposed by FCVG [65] or methods like AnyCharV [62] which aim for broader character controllability.
@@ 4.1.2 Depth-Guided Generation
Several works have explored depth-guided video synthesis. ControlVideo [88] and Control-A-Video [87] demonstrate how various control signals, including depth, could be integrated into large pre-trained text-to-image diffusion models to guide video generation. SparseCtrl [84] focuses on using sparse depth controls. DreamDance [82] leverages depth for dynamic 3D scene generation. Methods like GD-VDM [85] explicitly incorporate geometric priors using depth. ControlNeXt [67] and Champ [69] also support depth guidance among other control types. Depth information is typically encoded and injected into the diffusion modelś U-Net architecture, often concatenated with noisy latents or fed into cross-attention layers. Some methods may also use depth-based loss functions to further encourage geometric fidelity, for instance, by comparing the depth map of a generated frame with the target depth map.
Challenges in depth-guided generation include the difficulty of obtaining accurate, dense, and temporally consistent depth maps, especially from monocular RGB videos. Errors or noise in the input depth maps can propagate to the generated video, leading to distorted geometry. Ensuring that the generated RGB frames are consistent with the conditioning depth information is non-trivial. Dynamic scenes with rapidly changing depths or complex non-rigid object deformations pose further difficulties. Moreover, incorporating 3D awareness often increases computational complexity. Approaches like Make-Your-Video [86] and Moonshot [83] aim to improve fidelity and consistency in such scenarios.
@@ 4.1.3 Landmark-Guided Generation
Landmark control signals can be provided as static or dynamic landmark sequences extracted from images or videos, or as high-level descriptions that are converted into landmark trajectories. For instance, TASTE-Rob [60] uses temporal landmark sequences to guide video generation. Follow-Your-Pose [2] and Follow-Your-Emoji [96] allow pose-based and emoji-driven landmark control, respectively, offering users intuitive ways to define movements and expressions.
Recent landmark-guided methods leverage landmarks to condition generative models, ensuring spatial alignment and temporal coherence. HunyuanPortrait [91] introduces implicit conditioning mechanisms to generate highly realistic and controllable portrait animations. Takin-ADA [94] and EchoMimic [4] specialize in audio-driven portrait animation, aligning facial landmarks with speech dynamics. EchoMimicV2 [93] extends this by simplifying body and facial control for efficient synthesis. LivePortrait [95] and IPT2V [92] emphasize identity preservation while animating portraits, ensuring that synthesized videos remain faithful to the input identity.
@@ 4.1.4 Sketch-Guided Generation
In previous works, SketchBetween [114] and Sketch Me A Video [115] both introduce a VAE-based framework for the sketch-based video generation application. Recently, the ControlNet structure [373] has been applied to sketch-guided video generation, but directly using it does not ensure the frame-to-frame consistency essential for coherent video production. For example, for enhancing the controllable video generation, STF [112] builds on the Text-to-Video Zero structure [374], combining it with the optimized latent codes, the modified attention mechanism, and the ControlNet structure. As the existing control method commonly relies on dense temporal maps, SparseCtrl [84] proposed an auxiliary encoder like ControlNet, which utilizes the sparse condition maps to generate the target video. Similarly, EasyControl [106] and CTRL-Adapter [107] designed a control adapter that supports various inputs, including image, depth, and sketch, facilitating the performance of controllable video generation applications. VidSketch [99] introduced an architecture for hand-drawn sketch-driven video generation, which uses the sketch control strength strategy and the TempSpatial attention mechanism, benefiting the different drawing level users and the video consistency. OpenSoraPlan [101] proposed a powerful video generation framework with a condition encoder.
Numerous works also focus on exploring sketch-guided video generation applications, such as animation, video colorization, and many other tasks. TaleCrafter [111], a visual storytelling system, designs a controllable text-to-image component that generates target images based on input conditions and converts them into video through a subsequent Image-to-Video (I2V) module. ToonCrafter [103] proposes a cartoon video interpolation method, which designs a sketch encoder to provide efficient drawing tools. AniDoc [100] and LVCD [102] also focus on the sketch-guided animation application. For general video generation, methods like MakePixelsDance [110], VideoLCM [109], FrameGuidance [89], OmniVDiff [90], and TF-T2V [108], these methods not only improve the quality and the controllability of the video but also optimize the consistency for the long-time video generation task. Besides, MagicStick [104], SketchVideo [113], MotionClone [105], and CameraCtrl [98] utilize the sketch-guided controllable component for video editing and camera motion video generation.
@@ 4.1.5 BBox-Guided Generation
To inject bbox information into generative models, researchers have developed various strategies, which can generally be categorized into two main types: cross-attention-based injection and additive encoder-based injection. MagicDrive [123], aiming for fine-grained control over 3D geometry, pioneers the separate handling of different layout types. For variable-length sequential data like 3D bounding boxes, it injects them via a cross-attention mechanism, similar to how text is processed. Delphi [120] also injects layout embeddings via cross-attention. Another early work, DriveDreamer [122], seeking to build a real-world driving model, fuses the representations of 3D bounding boxes ($B_{i}\in\mathbb{R}^{N\times N_{B}\times 16}$ ) with visual features using a gated self-attention mechanism. To tackle the challenge of learning box-object correlations from purely visual cues, Boximator [125] introduces a new self-attention layer into the existing U-Net blocks, which processes control tokens derived from Fourier-encoded box coordinates and object IDs. Panacea [127] adopts a different strategy to generate panoramic videos. It first projects all 3D layouts (including boxes and road maps) into the camera’s perspective, converting them into a multi-channel control map, and then uses the ControlNet architecture for unified control. Building on these works, DreamForge [118] further extends the application of ControlNet to achieve more precise geometric control and longer video generation. It not only introduces Perspective Guidance to help the network generate geometrically accurate scenes but also designs Object-wise Position Encoding (OPE) to enhance foreground object modeling by encoding frustum sampling points within 3D bounding boxes. Ctrl-V [126] also utilizes a ControlNet architecture but introduces a unique two-stage pipeline. To achieve high-fidelity control from only start and end box positions, it first employs a diffusion-based BBox Generator to predict a complete video of the bounding box trajectories. These trajectory videos, which render boxes directly into pixel space, then serve as the conditional input for a Box2Video generation network.
@@
Moreover, some methods explore different injection methods. LLM-Grounded Video Diffusion (LVD) [121] takes a novel approach by leveraging a Large Language Model (LLM) to generate Dynamic Scene Layouts from text prompts. Then, during inference, it guides the attention maps using an energy function, $E_{topk}=-\text{Topk}(A\cdot M)+\text{Topk}(A\cdot(1-M))$, thereby enhancing the model’s ability to adhere to complex spatiotemporal dynamics without additional training. Also operating without training, TrailBlazer [222] directly edits spatial and temporal attention maps during the initial denoising steps to guide subjects along key-framed bounding box trajectories. Recently, architectures based on the DiT have gained attention for their excellent scalability. For instance, DIVE [119], aiming to generate multi-view consistent videos, integrates a ControlNet-Transformer into the DiT architecture to process road information and uses a joint cross-attention mechanism to fuse scene and instance layouts. MagicDrive-DiT [124] focuses on solving the mismatch between the spatio-temporal latent codes from a 3D VAE and per-frame geometric control signals. It designs a novel spatial-temporal conditional encoding module to ensure precise geometric control for high-resolution long videos. Meanwhile, DriveScape [117], to achieve effective control from sparse conditions, proposes a Bi-Directional Modulated Transformer that aligns and fuses multi-modal information through a series of latent-to-condition and condition-to-image attention steps. To facilitate its primary goal of jointly generating consistent 2D-3D multi-modal data, HoloDrive [116] employs a T2I-Adapter-like structure to flexibly inject projected 3D box and map conditions.
In contrast to the aforementioned diffusion-based methods, some works leverage token-based transformers. For instance, FACTOR [375] is a non-diffusion model that operates on discrete video tokens, employing a masked bidirectional transformer to generate video, enabling fine-grained control by injecting sparse, user-drawn box trajectories and reference images through newly introduced adaptive cross-attention layers.

### v2

@@ 5.1.1 Pose-Guided Generation
Pose-guided video generation specifically aims to animate subjects, predominantly humans, according to a sequence of predefined poses. This is a critical task for applications such as virtual avatar animation, character generation for interactive entertainment, human motion synthesis from sparse inputs, and fashion video synthesis [karras2023dreamposefashionimagetovideosynthesis, jin2024alignment, wang2025diffdecompose, gong2025relationadapter, ma2024magicstickcontrollablevideoediting, ma2025followcreation, wang2024cove]. The core challenge lies in generating temporally coherent video frames where the subject’s appearance is preserved while accurately following the target pose sequence $P={p_{1},p_{2},...,p_{T}}$, where each $p_{t}$ represents the pose in frame $t$.
@@
The input pose information $p_{t}$ is typically represented as 2D skeletal keypoints (e.g., COCO format), 3D skeletal coordinates, or more dense representations like DensePose [guler2018densepose], which maps image pixels to 3D surface coordinates of the human body. These pose sequences can be extracted from existing videos using off-the-shelf pose estimators (e.g., OpenPose [cao2019openpose]), derived from motion capture (MoCap) data, or generated algorithmically. The choice of pose representation influences the granularity of control and the complexity of the generation task.
@@
Numerous approaches have been proposed to achieve pose-guided generation. Early methods [yang2018poseguidedhumanvideo] often rely on GANs, using pose information to condition the generator, for instance, by rendering pose stick figures and feeding them as input alongside a latent code. More recent work leverages the power of diffusion models. For example, MagicAnimate [xu2023magicanimatetemporallyconsistenthuman] and Animate Anyone [hu2024animateanyoneconsistentcontrollable] and its successor Animate Anyone 2 [hu2025animate2highfidelitycharacter] utilize diffusion models conditioned on pose sequences and a reference image to generate temporally consistent animations. Techniques like Champ [zhu2024champ] and MimicMotion [zhang2024mimicmotionhighqualityhumanmotion] focus on high-fidelity human motion synthesis. DirectorLLM [song2024directorllmhumancentricvideogeneration] employs large language models to direct human-centric video generation based on pose and textual descriptions. Follow Your Pose [ma2024followposeposeguidedtexttovideo] introduces a two-stage training strategy to create high-quality character videos. ControlNet-based approaches [zhang2023addingconditionalcontroltexttoimage] have also been adapted, where pose maps serve as direct spatial conditions for pre-trained text-to-image diffusion models, which are then fine-tuned or augmented with temporal modules for video tasks [peng2024controlnext, wang2024discodisentangledcontrolrealistic]. Disentangling pose from appearance is a key strategy, often achieved by dedicated appearance encoders and pose encoders, with mechanisms like feature warping or attention to align appearance features with the target pose [li2025disposedisentanglingposeguidance]. Temporal consistency is often enforced through temporal attention mechanisms across frames or by incorporating temporal smoothness losses.
Table 2: Representative works of controllable generation with video generative model.
Type Method Venue Model Condition Training Dataset
Structure Control Follow-Your-Pose [ma2024followposeposeguidedtexttovideo] AAAI 2024 UNet Pose, Landmark, Text LAION-400M [schuhmann2021laion-400m], HDVILA [xue2021hdvila]
Make-Your-Video [xing2024make] TVCG 2024 UNet Depth, Text WebVid-10M [Bain21WebVid-10M]
ToonCrafter [10.1145/3687761] TOG 2024 UNet sketch, Image, Text Self-Construction Datasets
DriveDreamer [zhao2025drivedreamer] AAAI 2024 UNet BBox, Image, Text nuScenes [caesar2019nuScenes]
EchoMimic [meng2024echomimicv2strikingsimplifiedsemibody] AAAI 2024 UNet Audio, Landmarks, Image HDTF [zhang2021HDTF], CelebV-HQ [zhu2022CelebV-HQ]
ID Control VideoBooth [jiang2023videobooth] CVPR 2023 UNet Object, Text WebVid-10M [Bain21WebVid-10M]
Vlogger [zhuang2024vlogger] CVPR 2024 UNet Object, Text WebVid-10M [Bain21WebVid-10M], LAION-400M [schuhmann2021laion-400m]
Phantom [liu2025phantom] ICCV 2025 DiT Subject, Text Panda-70M [chen2024panda-70m], Subject200k [tan2024subject200k], OmniGen [xiao2024OmniGen]
Image Control VideoCrafter1 [chen2023videocrafter1opendiffusionmodels] arXiv 2023 UNet Image, Text LAION-COCO-600M [LAION-COCO], WebVid-10M [Bain21WebVid-10M]
Lumiere [10.1145/3680528.3687614] SIGGRAPH-Asia 2024 UNet Image, Mask, Style, Text Self-Construction Datasets
ConsistI2V [ren2024consisti2venhancingvisualconsistency] TMLR 2024 UNet Image, Text WebVid-10M [Bain21WebVid-10M]
Follow-Your-Click [ma2024followyourclickopendomainregionalimage] AAAI 2025 UNet Click, Image, Text WebVid-10M [Bain21WebVid-10M]
NOVA [deng2025autoregressivevideogenerationvector] ICLR 2025 Autoregressive Image, Text Panda-70M [chen2024panda-70m], Pexels [Pexels]
Temporal Control Motion-I2V [10.1145/3641519.3657497] SIGGRAPH 2024 UNet Trajectory, Image, Text WebVid-10M [Bain21WebVid-10M]
MotionBooth [wu2024motionbooth] NeurIPS 2024 UNet Motion, Camera, BBox, Image, Text Panda-70M [chen2024panda-70m]
ViewCrafter [wu2024motionbooth] TPAMI 2025 UNet Camera, Image RealEstate10K [tinghui2018RealEstate10K], DL3DV [ling2023DL3DV]
Direct-A-Video [yang2024directavideo] SIGGRAPH 2024 UNet Camera, Text MovieShot [rao2020MovieShot]
Audio Control DAA2V [yariv2024diverse] AAAI 2023 UNet Audio, Image VGGSound [Chen20VGGSound], Landscape [lee2022Landscape], AudioSet-Drums [gemmeke2017AudioSet]
MOFA-Video [niu2024mofavideo] ECCV 2024 UNet Trajectory, Audio, Flow, Image WebVid-10M [Bain21WebVid-10M]
EMO [tian2024emoemoteportraitalive] ECCV 2024 UNet Audio, Image HDTF [zhang2021HDTF], VFHQ [xie2022VFHQ], CelebV-HQ [zhu2022CelebV-HQ]
MotionCraft [bian2025motioncraft] AAAI 2025 DiT Music, Speech, Text HumanML3D [Guo_2022_CVPR_HumanML3D], BEAT2 [liu2023BEAT2], FineDance [li2023finedance]
Other Control Panacea [wen2024panacea] CVPR 2023 UNet BEV, Text nuScenes [caesar2019nuScenes]
StyleMaster [ye2024stylemaster] CVPR 2025 DiT Style, Video, Text Self-Construction Datasets
UniVST [song2024univst] TPAMI 2025 DiT Style, Mask, Video None
GS-DiT [bian2025gs] CVPR 2025 DiT Point, Video, Text WebVid-10M [Bain21WebVid-10M]
Universal Control VideoComposer [wang2023videocomposer] NeurIPS 2023 UNet Universal Conditions WebVid-10M [Bain21WebVid-10M], LAION-400M [schuhmann2021laion-400m]
FullDiT [ju2025fulldit] ICCV 2025 DiT Universal Conditions MiraData [ju2024miradata], RealEstate10K [tinghui2018RealEstate10K], ConceptMaster [huang2025conceptmaster], Panda-70M [chen2024panda-70m]
VACE [jiang2025vace] ICCV 2025 DiT Universal Conditions Self-Construction Datasets
@@
Despite significant progress, pose-guided generation still faces challenges. Handling severe self-occlusions, where parts of the subject are hidden due to the pose, requires robust inpainting capabilities. Maintaining fine-grained texture details and consistent identity across large pose variations remains difficult, especially when appearance must be reliably transferred from a reference image without explicit identity control. Ensuring natural and smooth transitions between poses, avoiding jittery or robotic movements, is crucial for realism. Furthermore, generalization to unseen poses, body shapes, or complex clothing items not well-represented in training data can be problematic, often requiring extensive and diverse datasets like those proposed by FCVG [zhu2024generativeinbetweeningframewiseconditionsdriven], which aims for broader character controllability.
@@ 5.1.2 Depth-Guided Generation
Several works have explored depth-guided video synthesis. ControlVideo [zhang2023controlvideo] and Control-A-Video [chen2023control] demonstrate how various control signals, including depth, could be integrated into large pre-trained text-to-image diffusion models to guide video generation. SparseCtrl [guo2024sparsectrl] focuses on using sparse depth controls. DreamDance [pang2024dreamdance] leverages depth for dynamic 3D scene generation. ControlNeXt [peng2024controlnext] and Champ [zhu2024champ] also support depth guidance among other control types. Depth information is typically encoded and injected into the diffusion modelś U-Net architecture, often concatenated with noisy latents or fed into cross-attention layers. Some methods may also use depth-based loss functions to further encourage geometric fidelity, for instance, by comparing the depth map of a generated frame with the target depth map.
Challenges in depth-guided generation include the difficulty of obtaining accurate, dense, and temporally consistent depth maps, especially from monocular RGB videos. Errors or noise in the input depth maps can propagate to the generated video, leading to distorted geometry. Ensuring that the generated RGB frames are consistent with the conditioning depth information is non-trivial. Dynamic scenes with rapidly changing depths or complex non-rigid object deformations pose further difficulties. Moreover, incorporating 3D awareness often increases computational complexity. Approaches like Make-Your-Video [xing2024make] and Moonshot [zhang2024moonshot] aim to improve fidelity and consistency in such scenarios.
@@ 5.1.3 Landmark-Guided Generation
Landmark control signals can be provided as static or dynamic landmark sequences extracted from images or videos, or as high-level descriptions that are converted into landmark trajectories. For instance, TASTE-Rob [zhao2025tasterobadvancingvideogeneration] uses temporal landmark sequences to guide video generation. Follow-Your-Pose [ma2024followposeposeguidedtexttovideo] and Follow-Your-Emoji [ma2024followyouremojifinecontrollableexpressivefreestyle] allow pose-based and emoji-driven landmark control, respectively, offering users intuitive ways to define movements and expressions.
Recent landmark-guided methods leverage landmarks to condition generative models, ensuring spatial alignment and temporal coherence. HunyuanPortrait [xu2025hunyuanportraitimplicitconditioncontrol] introduces implicit conditioning mechanisms to generate highly realistic and controllable portrait animations. Takin-ADA [lin2024takinadaemotioncontrollableaudiodriven] and EchoMimic [chen2024echomimiclifelikeaudiodrivenportrait] specialize in audio-driven portrait animation, aligning facial landmarks with speech dynamics. EchoMimicV2 [meng2024echomimicv2strikingsimplifiedsemibody] extends this by simplifying body and facial control for efficient synthesis. LivePortrait [guo2025liveportraitefficientportraitanimation] and IPT2V [yuan2024identitypreserving] emphasize identity preservation while animating portraits, ensuring that synthesized videos remain faithful to the input identity.
@@ 5.1.4 Sketch-Guided Generation
In previous works, works like SketchBetween [Loftsdottir_2022] introduce a VAE-based framework for the sketch-based video generation application. Recently, the ControlNet structure [zhang2023addingconditionalcontroltexttoimage] has been applied to sketch-guided video generation, but directly using it does not ensure the frame-to-frame consistency essential for coherent video production. As the existing control method commonly relies on dense temporal maps, SparseCtrl [guo2024sparsectrl] proposed an auxiliary encoder like ControlNet, which utilizes the sparse condition maps to generate the target video. Similarly, CTRL-Adapter [lin2024ctrladapterefficientversatileframework] designed a control adapter that supports various inputs, including image, depth, and sketch, facilitating the performance of controllable video generation applications. OpenSoraPlan [lin2024opensoraplanopensourcelarge] proposed a powerful video generation framework with a condition encoder.
Numerous works also focus on exploring sketch-guided video generation applications, such as animation, video colorization, and many other tasks. TaleCrafter [gong2023talecrafter], a visual storytelling system, designs a controllable text-to-image component that generates target images based on input conditions and converts them into video through a subsequent Image-to-Video (I2V) module. ToonCrafter [10.1145/3687761] proposes a cartoon video interpolation method, which designs a sketch encoder to provide efficient drawing tools. AniDoc [meng2025anidocanimationcreationeasier] and LVCD [Huang_2024] also focus on the sketch-guided animation application. For general video generation, methods like MakePixelsDance [zeng2023makepixelsdancehighdynamic], VideoLCM [wang2023videolcmvideolatentconsistency], and TF-T2V [Wang_2024_CVPR], these methods not only improve the quality and the controllability of the video but also optimize the consistency for the long-time video generation task. Besides, MagicStick [ma2024magicstickcontrollablevideoediting], SketchVideo [Liu_2025_CVPR], MotionClone [ling2024motionclonetrainingfreemotioncloning], and CameraCtrl [he2024cameractrl] utilize the sketch-guided controllable component for video editing and camera motion video generation.
Table 3: Quantitative comparison of Landmark-guided video generation on $512\times 512$ test images. LMD multiplied by $10^{-3}$, AED multiplied by $10^{-2}$, APD multiplied by $10^{-3}$ and Identity Similarity multiplied by $10^{-1}$. $\uparrow$} indicates higher is better. $\downarrow$} indicates lower is better. The best results are in bold.
Method Self Reenactment Cross Reenactment
LMD $\downarrow$ FID-VID $\downarrow$ FVD $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ AED $\downarrow$ APD $\downarrow$ Identity $\uparrow$ Facial $\uparrow$ Video $\uparrow$ Temporal $\uparrow$
Similarity Movements Quality Smoothness
LivePortrait [guo2025liveportraitefficientportraitanimation] 9.14 82.71 483.38 31.41 0.72 0.22 26.83 26.97 8.71 3.00 4.11 4.06
AniPortrait [wei2024aniportraitaudiodrivensynthesisphotorealistic] 6.67 81.90 430.24 30.54 0.67 0.27 22.42 21.32 7.95 3.83 4.00 3.81
FollowYE [ma2024followyouremojifinecontrollableexpressivefreestyle] 5.63 77.15 417.53 30.84 0.68 0.25 20.06 20.95 8.18 4.11 4.23 3.33
X-Potrait [xie2024x] 6.23 82.93 416.41 30.81 0.71 0.19 20.66 20.39 8.03 3.88 3.73 3.49
HunyuanPortrait [xu2025hunyuanportraitimplicitconditioncontrol] 2.02 75.81 333.48 32.98 0.81 0.11 19.45 19.20 8.87 4.55 4.69 4.61
@@ 5.1.5 BBox-Guided Generation
To inject bbox information into generative models, researchers have developed various strategies, which can generally be categorized into two main types: cross-attention-based injection and additive encoder-based injection. MagicDrive [gao2023magicdrive], aiming for fine-grained control over 3D geometry, pioneers the separate handling of different layout types. Delphi [ma2024unleashing] also injects layout embeddings via cross-attention. Another early work, DriveDreamer [zhao2025drivedreamer], seeking to build a real-world driving model, fuses the representations of 3D bounding boxes ($B_{i}\in\mathbb{R}^{N\times N_{B}\times 16}$ ) with visual features using a gated self-attention mechanism. To tackle the challenge of learning box-object correlations from purely visual cues, Boximator [wang2024boximator] and Panacea [wen2024panacea] introduce the new self-attention layer into the existing U-Net blocks, which processes control tokens derived from Fourier-encoded box coordinates and object IDs. Ctrl-V [luo2025ctrl] also utilizes a ControlNet architecture but introduces a unique two-stage pipeline. It first employs a diffusion-based BBox Generator to predict a complete video of the bounding box trajectories.
@@
Moreover, some methods explore different injection methods. LLM-Grounded Video Diffusion (LVD) [lian2023llm] takes a novel approach by leveraging a Large Language Model (LLM) to generate Dynamic Scene Layouts from text prompts. Also operating without training, TrailBlazer [ma2024trailblazer] directly edits spatial and temporal attention maps during the initial denoising steps to guide subjects along key-framed bounding box trajectories. Recently, architectures based on the DiT have gained attention for their excellent scalability. For instance, DIVE [jiang2024dive], aiming to generate multi-view consistent videos, integrates a ControlNet-Transformer into the DiT architecture to process road information and uses a joint cross-attention mechanism to fuse scene and instance layouts.
5.1.6 Summary
In video generation, structural control acts as spatial cues, providing the foundational framework for dynamic visuals. Its core function is to precisely translate various abstract spatial requirements—from precise joint poses to casual user sketches—into natural and plausible visual movements. Early GAN-based frameworks [yang2018poseguidedhumanvideo] often conditioned the generation process by directly rendering structural skeletons together with latent features. And this paradigm was subsequently dominated by UNet-based architectures which employs Adapter interfaces (e.g., ControlNet [zhang2023addingconditionalcontroltexttoimage])to graft spatial constraints onto main network. Currently, the research focus has clearly shifted towards diffusion Transformers(e.g., MagicDrive-DiT [gao2024magicdrivedit])due to the scaling limits of CNNs.
This shift is from the user’s need for finer control granularity and temporal coherence, with trends leaning toward integrated approaches that enable combining multiple structural signals (e.g., pose with depth in DreamDance [pang2024dreamdance]). Beyond interaction, profound technical challenges persist in physical awareness and generalization. Current methods often prioritize visual alignment over kinematic plausibility, leading to physically impossible deformations, and remain limited to specific domains (e.g., humans). Consequently, future improvement is needed for interaction simplicity and physics-aware world modeling. The ultimate goal is to bridge the gap between abstract control and concrete realization, ensuring outcomes are not only physically consistent but also faithfully aligned with the user’s creative intent.
Table 4: Quantitative comparison of landmark-guided video generation on TED-talks Dataset. L1 multiplied by $10^{-4}$. $\uparrow$} indicates higher is better. $\downarrow$} indicates lower is better. The best and secondary results are in bold and underline, respectively. MRAA* and TPSMM* proposed these methods using real videos as driving signals. Therefore, they are not involved in quantitative comparison.
Method FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ L1 $\downarrow$ FID-VID $\downarrow$ FVD $\downarrow$
MRAA*[siarohin2021motion] 25.76 0.807 32.62 0.216 0.36 12.20 126.28
TPSMM*[zhao2022thin] 24.69 0.819 32.92 0.205 0.35 9.04 87.76
MRAA [siarohin2021motion] 50.36 0.762 31.90 0.266 0.50 82.79 493.02
TPSMM [zhao2022thin] 23.71 0.771 32.30 0.252 0.49 32.12 260.67
DisCo [wang2024discodisentangledcontrolrealistic] 75.48 0.575 27.99 0.309 1.21 66.18 393.04
MagicAnimate [xu2023magicanimatetemporallyconsistenthuman] 41.58 0.529 28.28 0.310 1.73 33.61 223.54
MagicPose [chang2024magicposerealistichumanposes] 23.39 0.723 30.08 0.236 0.81 27.53 214.23
FollowYourPosev2 [xue2024follow] 18.21 0.779 30.88 0.198 0.46 10.24 81.73